Show that the binomial is a factor of the polynomial. Then factor the polynomial completely.
s(x)=x4+4x3−64x−256; x+4
s(x)=

Answer by stanbon(75887)

Show that the binomial is a factor of the polynomial. Then factor the polynomial completely.
s(x)=x^4+4x^3−64x−256; x+4
------
Using synthetic division::
-4)....1....4....0....-64....-256
.......1....0....0....-64...|..0
------------------------------------
s(x) = (x+4)(x^3-64)
------
s(x) = (x+4)(x-4)(x^2 + 4x + 16)
=======================
Cheers,
Stan H.
------------------


s(x)=